# Gaussian Process-Based Prediction and Control of Hammerstein-Wiener Systems

### v1

@@ V Numerical Examples
A sinusoidal output reference with lower and upper bound constraints is considered. No input constraints are enforced, i.e., $\mathcal{U}_{k}=\mathbb{R}^{L^{\prime}}$. The following parameters are used: $M=2$, $p=0.7$, $Q=R=\mathbb{I}$. To avoid infeasibilities, the tightened output constraints are implemented as soft constraints with a penalty weight of 100.
The closed-loop output trajectories are plotted in Figure 4. The results demonstrate that only the proposed control algorithm (37) is able to track the reference closely within the output bounds. The black-box Gaussian process model cannot provide any reasonable control due to its large posterior standard deviation leading to severe constraint violation throughout the simulation. The prediction performance of the proposed algorithm is further analyzed in Figure 5. It can be seen that the output predictions are very close to the true output trajectory and the posterior standard deviation provides a good estimate of the prediction error with the true trajectory staying within the one standard deviation range from the predictions for most of the time.
@@
Figure 4: Closed-loop output trajectories with different data-driven predictive controllers.
Figure 5: Prediction accuracy and error bounds of the proposed control algorithm.
VI Conclusions
@@
This work presents data-driven prediction and control algorithms of Hammerstein-Wiener systems by implicit Gaussian process regression. By formulating the predictor as an implicit function, a Gaussian process model is obtained by deriving kernel functions with Gaussian process priors for nonlinearities. A stable spline hyperprior is used to estimate linear model parameters as hyperparameters. Applied to receding horizon control, the algorithm minimizes the expected control cost and complies with output chance constraints. Validation through numerical examples demonstrates improved performance over the black-box Gaussian process method.
@@
One drawback of the proposed approach is its high computational complexity due to a high number of hyperparameters and complex kernel design. Future work may discuss optimization algorithms for solving the hyperparameter tuning problem. Closed-loop guarantees for the receding horizon control algorithm can also be investigated.

### v2

@@ 5.2 Data-Driven Predictive Control
The output nonlinearity is monotonically increasing and satisfies Assumption 2 with $M=0.5$. Unit Gaussian inputs are used to obtain the training data. A sinusoidal output reference with lower and upper bound constraints $0.2\leq\mathbf{y}^{k}-\mathbf{r}^{k}\leq 0.2$ is considered. No input constraints are enforced, i.e., $\mathcal{U}_{k}=\mathbb{R}^{L^{\prime}}$. The following parameters are used: $M=0.5$, $p=0.7$, $Q=100\cdot\mathbb{I}$, $R=\mathbb{I}$.
We have also implemented receding horizon control for the black-box GP and the linear predictors. For the black-box GP models, the following optimization problem is solved at time $k$:
$\underset{\mathbf{u}^{k},\mathbf{y}^{k}}{\mathrm{min}}$ $\ \left\lVert\mathbf{u}^{k}\right\rVert_{R}^{2}+\left\lVert\mathbf{y}^{k}-\mathbf{r}^{k}\right\rVert_{Q}^{2}+\mathrm{diag}\left(Q\right)^{\top}\boldsymbol{\sigma}^{2}_{\mathrm{bb}}(\xi^{k})$
$\mathrm{s.t.}$ $\ \mathbf{y}^{k}=\mathbf{m}_{\mathrm{bb}}(\xi^{k}),\ \mathbf{u}^{k}\in\mathcal{U}_{k},\$
$\ H\mathbf{y}^{k}\leq q-\sqrt{\mu_{\chi^{2}}(p)\mathrm{diag}\left(H\Sigma_{\mathrm{bb}}(\xi^{k})H^{\top}\right)},$
where $\xi^{k}=\mathrm{col}\left(u_{k-L_{0}},\dots,u_{k-1},\mathbf{u}^{k},y_{k-L_{0}},\dots,y_{k-1}\right)$ and $\Sigma_{\mathrm{bb}}(\xi^{k})$ denotes the diagonal matrix constructed from $\boldsymbol{\sigma}_{\mathrm{bb}}^{2}(\xi^{k})$. The derivation is similar to Section 4. To quantify the multi-step prediction error for the recursive GP model, we tried to apply the linear approximation approach [15], but this leads to infeasible problems at all times due to too conservative error quantification. Instead, we ignore the prediction error due to noisy regression inputs. For the linear predictor, we apply subspace predictive control [13] with certainty equivalence. In addition, we have applied nonlinear MPC using the true nonlinear model as a benchmark for ideal performance.
The closed-loop input-output trajectories are plotted in Figure 5. The results demonstrate that the proposed control algorithm (56) achieves the best performance among the data-driven algorithms. It performs very similarly to the nonlinear MPC algorithm with the true model and is able to track the reference closely within the output bounds. The black-box GP algorithms and the linear predictor under-actuate at the peaks of the output reference, indicating that they are not accurately capturing the output nonlinearity of the model.
Similar to prediction, our proposed method requires a longer computation time. The median computation time per iteration is 4.41 s, 0.24 s, 0.13 s, 0.30 s, and 0.01 s for the DDPC algorithm (56), multi-step black-box GP, recursive black-box GP, subspace predictive control, and the nonlinear MPC, respectively.
@@
Figure 5: Closed-loop input-output trajectories with different data-driven predictive controllers.
6 Conclusions
@@
This work presents data-driven prediction and control algorithms of Hammerstein-Wiener systems by implicit GP regression. By formulating the predictor as an implicit function, a Gaussian process (GP) model is obtained by deriving kernel functions with linear model parameters and GP priors for nonlinearities. A stable spline hyperprior is used to estimate linear model parameters as hyperparameters. Monotonicity information of the nonlinearities can be included by expectation propagation with virtual derivative points. Applied to receding horizon control, the algorithm minimizes the expected control cost and complies with output chance constraints. Validation through numerical examples demonstrates improved performance over black-box GP methods.
@@
One drawback of the proposed approach is its high computational complexity due to a high number of hyperparameters and complex kernel design. Future work may discuss tailored algorithms for solving the optimization problems. Closed-loop guarantees for the receding horizon control algorithm can also be investigated.
Appendix A Derivation of (22)-(24)
Let
$\mathbf{u}_{k}^{d}=\mathrm{col}\left(u^{d}_{k},\dots,u^{d}_{k+L-1}\right),\ \mathbf{y}_{k}^{d}=\mathrm{col}\left(y^{d}_{k},\dots,y^{d}_{k+L-1}\right),$
such that $\mathrm{vec}\left(H_{u}\right)=\mathrm{col}\left(\mathbf{u}_{1}^{d},\dots,\mathbf{u}_{M}^{d}\right)$ and $\mathrm{vec}\left(H_{y}\right)=\mathrm{col}\left(\mathbf{y}_{1}^{d},\dots,\mathbf{y}_{M}^{d}\right)$. From (18) and (20), we have
$m(\eta_{k}^{d})=\left[\Gamma_{1}\ \bar{\Gamma}_{2}\right]\mathrm{col}\left(\mathbf{m}_{u}(\mathbf{u}_{k}^{d}),\mathbf{m}_{y}(\mathbf{y}_{k}^{d})\right).$
This leads to
$\left[m(\eta^{d}_{1})\ \cdots\ m(\eta^{d}_{M})\right]=\left[\Gamma_{1}\ \bar{\Gamma}_{2}\right]\begin{bmatrix}\mathbf{m}_{u}(H_{u})\\ \mathbf{m}_{y}(H_{y})\end{bmatrix}$.
Vectorizing both sides leads to (22).
From (18) and (21), we have
$k(\eta_{k}^{d},\eta)=\Gamma_{1}\mathbf{k}_{u}(\mathbf{u}_{k}^{d},\mathbf{u})\Gamma_{1}^{\top}+\bar{\Gamma}_{2}\mathbf{k}_{y}(\mathbf{y}_{k}^{d},\mathbf{y})\bar{\Gamma}_{2}^{\top}.$
This leads to
$\begin{aligned} &\mathbf{k}(\boldsymbol{\eta}^{d},\eta)=\mathrm{col}\left(k(\eta^{d}_{1},\eta),\dots,k(\eta^{d}_{M},\eta)\right)\\ =&\left(\mathbb{I}\otimes\Gamma_{1}\right)\begin{bmatrix}\mathbf{k}_{u}(\mathbf{u}_{1}^{d},\mathbf{u})\\ \vdots\\ \mathbf{k}_{u}(\mathbf{u}_{M}^{d},\mathbf{u})\end{bmatrix}\Gamma_{1}^{\top}+\left(\mathbb{I}\otimes\bar{\Gamma}_{2}\right)\begin{bmatrix}\mathbf{k}_{y}(\mathbf{y}_{1}^{d},\mathbf{y})\\ \vdots\\ \mathbf{k}_{y}(\mathbf{y}_{M}^{d},\mathbf{y})\end{bmatrix}\bar{\Gamma}_{2}^{\top}\\ =&\left(\mathbb{I}\otimes\Gamma_{1}\right)\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathbf{u})\Gamma_{1}^{\top}+\left(\mathbb{I}\otimes\bar{\Gamma}_{2}\right)\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathbf{y})\bar{\Gamma}_{2}^{\top}.\end{aligned}$
Similarly,
$k(\eta_{k}^{d},\eta_{k}^{d})=\Gamma_{1}\mathbf{k}_{u}(\mathbf{u}_{k}^{d},\mathbf{u}_{k}^{d})\Gamma_{1}^{\top}+\bar{\Gamma}_{2}\mathbf{k}_{y}(\mathbf{y}_{k}^{d},\mathbf{y}_{k}^{d})\bar{\Gamma}_{2}^{\top},$
and
$K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d})$
$=$ $\left(\mathbb{I}\otimes\Gamma_{1}\right)\begin{bmatrix}\mathbf{k}_{u}(\mathbf{u}_{1}^{d},\mathbf{u}_{1}^{d})&\cdots&\mathbf{k}_{u}(\mathbf{u}_{1}^{d},\mathbf{u}_{M}^{d})\\ \vdots&\ddots&\vdots\\ \mathbf{k}_{u}(\mathbf{u}_{M}^{d},\mathbf{u}_{1}^{d})&\cdots&\mathbf{k}_{u}(\mathbf{u}_{M}^{d},\mathbf{u}_{M}^{d})\end{bmatrix}\left(\mathbb{I}\otimes\Gamma_{1}^{\top}\right)$
$+$ $\left(\mathbb{I}\otimes\bar{\Gamma}_{2}\right)\begin{bmatrix}\mathbf{k}_{y}(\mathbf{y}_{1}^{d},\mathbf{y}_{1}^{d})&\cdots&\mathbf{k}_{y}(\mathbf{y}_{1}^{d},\mathbf{y}_{M}^{d})\\ \vdots&\ddots&\vdots\\ \mathbf{k}_{y}(\mathbf{y}_{M}^{d},\mathbf{y}_{1}^{d})&\cdots&\mathbf{k}_{y}(\mathbf{y}_{M}^{d},\mathbf{y}_{M}^{d})\end{bmatrix}\left(\mathbb{I}\otimes\bar{\Gamma}_{2}^{\top}\right)$
$=$ $\left(\mathbb{I}\otimes\Gamma_{1}\right)\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathrm{vec}\left(H_{u}\right))\left(\mathbb{I}\otimes\Gamma_{1}^{\top}\right)$
$\qquad+\left(\mathbb{I}\otimes\bar{\Gamma}_{2}\right)\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathrm{vec}\left(H_{y}\right))\left(\mathbb{I}\otimes\bar{\Gamma}_{2}^{\top}\right).$
Appendix B Details of the expectation propagation algorithm
Let $F_{\mathcal{N}}\left(\phi^{\prime}(y_{\mathrm{m},i})/\nu_{0}\right)$ be approximated by $t_{i}(\tilde{Z}_{i},\tilde{\mu}_{i},\tilde{\sigma}_{i}^{2}):=\tilde{Z}_{i}p_{\mathcal{N}(\tilde{\mu}_{i},\tilde{\sigma}_{i}^{2})}(\phi^{\prime}(y_{\mathrm{m},i}))$, where $p_{\mathcal{N}(\tilde{\mu}_{i},\tilde{\sigma}_{i}^{2})}(\phi^{\prime}(y_{\mathrm{m},i}))$ denotes the probability density function of $\mathcal{N}(\tilde{\mu}_{i},\tilde{\sigma}_{i}^{2})$. Let
$\tilde{\boldsymbol{\mu}}_{\mathrm{j}}$ $=\mathrm{col}\left(\tilde{\mu}_{1},\dots,\tilde{\mu}_{n_{m}},\boldsymbol{\chi}^{d}=\mathbf{0}\right),$ (57)
$\tilde{\Sigma}_{\mathrm{j}}$ $=\mathrm{blkdiag}\left(\tilde{\sigma}_{1}^{2},\dots,\tilde{\sigma}_{n_{m}}^{2},\bar{\Sigma}\right),$
$\mathbf{z}$ $=\mathrm{col}\left(\Phi^{\prime}(\mathbf{y}_{\mathrm{m}}),\mathbf{f}(\boldsymbol{\eta}^{d})\right).$
Define the approximate posterior distribution:
$q(\mathbf{z}|\mathcal{M},\boldsymbol{\chi}^{d}):=\frac{1}{Z}p(\mathbf{z})p(\boldsymbol{\chi}^{d}|\mathbf{f}(\boldsymbol{\eta}^{d}))\prod_{i=1}^{n_{m}}t_{i}(\tilde{Z}_{i},\tilde{\mu}_{i},\tilde{\sigma}_{i}^{2}),$ (58)
where $Z$ is the normalization term and $q(\cdot)$ denotes approximate probability densities. We would like to find parameters $\tilde{Z}_{i}$, $\tilde{\mu}_{i}$, and $\tilde{\sigma}_{i}^{2}$ such that
$q(\mathbf{z}|\mathcal{M},\boldsymbol{\chi}^{d})\approx p(\mathbf{z}|\mathcal{M},\boldsymbol{\chi}^{d}).$
Note that
$p(\boldsymbol{\chi}^{d}|\mathbf{f}(\boldsymbol{\eta}^{d}))\prod_{i=1}^{n_{m}}t_{i}(\tilde{Z}_{i},\tilde{\mu}_{i},\tilde{\sigma}_{i}^{2})=p_{\mathcal{N}(\tilde{\boldsymbol{\mu}}_{\mathrm{j}},\tilde{\Sigma}_{\mathrm{j}})}(\mathbf{z}).$
Then, from (29) and the Gaussian product rule, we have approximately
$\mathbf{z}|\mathcal{M},\boldsymbol{\chi}^{d}\sim\mathcal{N}\left(\boldsymbol{\mu},\mathbf{S}\right),$ (59)
where
$\boldsymbol{\mu}=\mathbf{S}\tilde{\Sigma}_{\mathrm{j}}^{-1}\tilde{\boldsymbol{\mu}}_{\mathrm{j}}+\mathbf{S}K_{\mathrm{j}}^{-1}\mathbf{m}_{\mathrm{j}},\quad\mathbf{S}=(K_{\mathrm{j}}^{-1}+\tilde{\Sigma}_{\mathrm{j}}^{-1})^{-1}.$
To achieve this, the EP algorithm starts from some initial approximate posterior and calculates the so-called cavity distribution for a particular virtual derivative point $y_{m,i}$. The cavity distribution is defined as the marginal posterior distribution $q(\phi^{\prime}(y_{\mathrm{m},i})|\mathcal{M},\boldsymbol{\chi}^{d})$ divided by the approximant $t_{i}(\tilde{Z}_{i},\tilde{\mu}_{i},\tilde{\sigma}_{i}^{2})$. Then we approximate the product of the cavity distribution and the likelihood $F_{\mathcal{N}}\left(\phi^{\prime}(y_{\mathrm{m},i})/\nu_{0}\right)$ with a Gaussian likelihood by matching their zero-th, first, and second moments. Finally, the approximant is updated such that the product of the cavity distribution and the approximant has the desired moments. This process iterates over all the virtual derivative points until convergence.
With the augmentation of the monotonicity information at the virtual derivative points, for any query point $\eta=\mathrm{col}\left(\mathbf{u},\mathbf{y}\right)$, we have approximately
[…]
where $\mathbf{k}_{\mathrm{j}}(\eta):=\mathrm{col}\left({\kappa^{\prime}_{y}}^{\top}(\eta,\mathbf{y}_{\mathrm{m}}),\mathbf{k}(\boldsymbol{\eta}^{d},\eta)\right)$. This leads to the conditional probability:
$f(\eta)|\mathbf{z}\sim\mathcal{N}\Big(m(\eta)+\mathbf{k}^{\top}_{\mathrm{j}}(\eta)K_{\mathrm{j}}^{-1}(\mathbf{z}-\mathbf{m}_{\mathrm{j}}),\\ k(\eta,\eta)-\mathbf{k}^{\top}_{\mathrm{j}}(\eta)K_{\mathrm{j}}^{-1}\mathbf{k}_{\mathrm{j}}(\eta)\Big).$ (61)
From the definition of conditional probability, we have
$q\left(f(\eta)|\mathcal{M},\boldsymbol{\chi}^{d}\right)=p\left(f(\eta)|\mathbf{z}\right)q(\mathbf{z}|\mathcal{M},\boldsymbol{\chi}^{d}).$
From (59) and (61), the posterior mean and covariance of $f(\eta)|\mathcal{M},\boldsymbol{\chi}^{d}$ are given by
$\mathbb{E}\left[f(\eta)|\mathcal{M},\boldsymbol{\chi}^{d}\right]=\int\mathbb{E}\left[f(\eta)|\mathbf{z}\right]q(\mathbf{z}|\mathcal{M},\boldsymbol{\chi}^{d})\mathrm{d}\mathbf{z}$ (62)
$=$ $m(\eta)-\mathbf{k}^{\top}_{\mathrm{j}}(\eta)K_{\mathrm{j}}^{-1}\mathbf{m}_{\mathrm{j}}+\mathbf{k}^{\top}_{\mathrm{j}}(\eta)K_{\mathrm{j}}^{-1}\boldsymbol{\mu}$
$=$ $m(\eta)+\mathbf{k}^{\top}_{\mathrm{j}}(\eta)(K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}})^{-1}\left(\tilde{\boldsymbol{\mu}}_{\mathrm{j}}-\mathbf{m}_{\mathrm{j}}\right),$
$\mathrm{cov}\left(f(\eta)|\mathcal{M},\boldsymbol{\chi}^{d}\right)$ (63)
$=$ $k(\eta,\eta)-\mathbf{k}^{\top}_{\mathrm{j}}(\eta)K_{\mathrm{j}}^{-1}\mathbf{k}_{\mathrm{j}}(\eta)+\mathbf{k}^{\top}_{\mathrm{j}}(\eta)K_{\mathrm{j}}^{-1}SK_{\mathrm{j}}^{-1}\mathbf{k}_{\mathrm{j}}(\eta)$
$=$ $k(\eta,\eta)-\mathbf{k}^{\top}_{\mathrm{j}}(\eta)(K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}})^{-1}\mathbf{k}_{\mathrm{j}}(\eta),$
where we used the property:
$\left(K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}}\right)^{-1}=K_{\mathrm{j}}^{-1}\left(K_{\mathrm{j}}^{-1}+\tilde{\Sigma}_{\mathrm{j}}^{-1}\right)^{-1}\tilde{\Sigma}_{\mathrm{j}}^{-1}.$